9. Refer to the graph of k below. 

(a) Evaluate k(2). 

(b) Write an equation for k(t). 

(c) Use your equation from part 9b to evaluate k(10).

(d) Use your equation from part 9b to solve k(t) = 16. Use Desmos to check your answer.
Answer #1

from graph K(2) when value of t=2 function is K(2) = 4 Vertex is (hik) = (2,4) passes through (1,7) y = a (xx-h7+k y = a (x-2k (i) - 16 364-2)५= 16 3 -2) - 12 4-2)-- t-2 = + 4 4-2 = +2 = 4, = 2 + 2 -2 + 2- =५ 0